# Generalization of Geometric Graph Neural Networks

### v1

@@ IV-B Cora Dataset
The Cora dataset is a smaller graph than the Arxic dataset. The task is also node label prediction, and in this case, there are $7$ classes. In this dataset, the input feature dimension is $1,433$, and there are $2,708$ nodes and $5,429$ edges [40]. In this case the experiments were done for $\{16,32,64\}$ hidden units and $\{2,3\}$ layers.
@@
Visually, we verify that the generalization gap follows a linear decrease in log scale with respect to the number of nodes in the training set. This gap is driven by the increase in the testing accuracy/loss given that the training value remains constant. Formally, we can see in Table I that the Pearson correlation coefficient is always above $0.96$ (except $2$ layers $64$ hidden units measured in the loss) which again shows a strong linear correlation between the number of nodes in the training set (in log scale) and the generalization gap.
@@ V Conclusion
In this paper, we implemented a manifold convolutional neural network as a limit model for graph neural networks. We considered the case when graphs are constructed from samples from the manifold. We showed that when a graph is constructed with the i.i.d. randomly sampled points over the manifold, the GNN converges to the MNN with a non-asymptotic rate both in probability and in expectation. We utilized these convergence results as a stepping stone towards proving an upper bound on the generalization gap of GNNs. Formally, we showed that the difference between the statistical and empirical risks decreases with the number of sampled points from the manifold while it increases with the dimension of the manifold. Furthermore, our bound admits a linear interpretation in a logarithmic scale which can be validated empirically. Through experiments, we verified our convergence and generalization results numerically with two real-world datasets. We trained GNNs on a varying number of nodes and compared their generalization gaps. We empirically verified that the theoretically proposed rate matches the one seen in practice for a variety of layers and hidden dimensions. We also computed the Pearson Correlation coefficient to validate the linear relationships and found values always above $0.9$, which indicates a strong linear relationship.
@@ A-A Proof of Proposition 1
[35, Theorem 2.4, Theorem 2.6] Let ${\mathcal{M}}\subset{\mathbb{R}}^{\mathsf{N}}$ of measure $\mu$ with probability density function $\rho$ be equipped with a weighted Laplace operator ${\mathcal{L}}_{\rho}$ as defined in (9), whose eigendecomposition is given by $\{\lambda_{i},\bm{\phi}_{i}\}_{i=1}^{\infty}$. Let ${\mathbf{L}}_{N}$ be the discrete graph Laplacian of graph weights defined as (15), with spectrum $\{\lambda_{i,N},\bm{\phi}_{i,N}\}_{i=1}^{N}$. Fix $K\in\mathbb{N}^{+}$ and assume that $\epsilon=\epsilon(N)\geq\left({\log(CN/\delta)}/{N}\right)^{1/(d+4)}$ Then, with probability at least $1-\delta$, we have
@@
$\leq\sum_{i=M+1}^{N}\left|\hat{h}(\lambda_{i,N})\right|\|{\mathbf{P}}_{N}f\|$ (46)
@@ A-C Proof of Theorem 1
$GA\leq R_{S}({\mathbf{H}}_{E})-R_{E}({\mathbf{H}}_{E}),$ (62)
$=\mathbb{E}_{X^{N}\sim\mu^{N}}\left[\ell\left(\bm{\Phi}({\mathbf{H}}_{E},{\mathbf{x}}_{N}),{\mathbf{y}}_{N}\right)\right]-\ell(\bm{\Phi}({\mathbf{H}}_{E},{\mathbf{x}}_{N}),{\mathbf{y}}_{N})$ (63)
We can now subtract and add $\ell(\bm{\Phi}({\mathbf{H}}_{E},{\mathcal{L}},f),g)$, as follows
$GA\leq\left(\mathbb{E}_{X^{N}\sim\mu^{N}}\left[\ell\left(\bm{\Phi}({\mathbf{H}}_{E},{\mathbf{x}}_{N}),{\mathbf{y}}_{N}\right)\right]-\ell(\bm{\Phi}({\mathbf{H}}_{E},{\mathcal{L}},f),g)\right)$
$\qquad\qquad+\left(\ell(\bm{\Phi}({\mathbf{H}}_{E},{\mathcal{L}},f),g)-\ell(\bm{\Phi}({\mathbf{H}}_{E},{\mathbf{x}}_{N}),{\mathbf{y}}_{N})\right)$ (64)
@@
$GA\leq\underbrace{\Bigg{|}\mathbb{E}_{X^{N}\sim\mu^{N}}\left[\ell\left(\bm{\Phi}({\mathbf{H}}_{E},{\mathbf{x}}_{N}),{\mathbf{y}}_{N}\right)\right]-\ell(\bm{\Phi}({\mathbf{H}}_{E},{\mathcal{L}},f),g)\Bigg{|}}_{\textbf{LHS}}$
$\qquad\qquad+\underbrace{\Bigg{|}\ell(\bm{\Phi}({\mathbf{H}}_{E},{\mathcal{L}},f),g)-\ell(\bm{\Phi}({\mathbf{H}}_{E},{\mathbf{x}}_{N}),{\mathbf{y}}_{N})\Bigg{|}}_{\textbf{RHS}}.$ (65)
We assume the loss function to be $L_{2}$ loss. The LHS in (65) can be decomposed as
$\left|\ell(\bm{\Phi}({\mathbf{H}}_{E},{\mathcal{L}},f),g)-\mathbb{E}_{X_{N}\sim\mu^{N}}\left[\ell\left(\bm{\Phi}({\mathbf{H}}_{E},{\mathbf{L}}_{N},{\mathbf{x}}_{N}),{\mathbf{y}}_{N}\right)\right]\right|$
$=\left|\|\bm{\Phi}({\mathbf{H}}_{E},{\mathcal{L}},f)-g\|_{\mathcal{M}}-\mathbb{E}_{X_{N}}[\|\bm{\Phi}({\mathbf{H}}_{E},{\mathbf{L}}_{N},{\mathbf{x}}_{N})-{\mathbf{y}}_{N}\|]\right|.$ (66)
By subtracting and adding an intermediate term $\mathbb{E}_{X_{N}}[{\mathbf{P}}_{N}\bm{\Phi}({\mathbf{H}}_{E},{\mathcal{L}},f)]$in the expectation value and the triangle inequality, we can decompose the expectation term as follows with ${\mathbf{y}}_{N}={\mathbf{P}}_{N}g$.
$\left|\mathbb{E}_{X_{N}}[\|\bm{\Phi}({\mathbf{H}}_{E},{\mathbf{L}}_{N},{\mathbf{x}}_{N})-{\mathbf{y}}_{N}\|]-\|\bm{\Phi}({\mathbf{H}}_{E},{\mathcal{L}},f)-g\|_{\mathcal{M}}\right|$
$\leq\Big{|}\mathbb{E}_{X_{N}}[\|\bm{\Phi}({\mathbf{H}}_{E},{\mathbf{L}}_{N},{\mathbf{x}}_{N})-{\mathbf{P}}_{N}\bm{\Phi}({\mathbf{H}}_{E},{\mathcal{L}},f)\|]+$
$\mathbb{E}_{X_{N}}[\|{\mathbf{P}}_{N}\bm{\Phi}({\mathbf{H}}_{E},{\mathcal{L}},f)-{\mathbf{P}}_{N}g\|]-\|\bm{\Phi}({\mathbf{H}}_{E},{\mathcal{L}},f)-g\|_{\mathcal{M}}\Big{|}$ (67)
$\leq\left|\mathbb{E}_{X_{N}}[\|\bm{\Phi}({\mathbf{H}}_{E},{\mathbf{L}}_{N},{\mathbf{x}}_{N})-{\mathbf{P}}_{N}\bm{\Phi}({\mathbf{H}}_{E},{\mathcal{L}},f)\|]\right|+$
$\left|\mathbb{E}_{X_{N}}[\|{\mathbf{P}}_{N}\bm{\Phi}({\mathbf{H}}_{E},{\mathcal{L}},f)-{\mathbf{P}}_{N}g\|]-\|\bm{\Phi}({\mathbf{H}}_{E},{\mathcal{L}},f)-g\|_{\mathcal{M}}\right|.$ (68)
The first term in (68) is the conclusion in Corollary 2, while the second term in (68) can be derived based on inequality 29, as we have
$\mathbb{P}\Bigg{(}\Big{|}\|{\mathbf{P}}_{N}(\bm{\Phi}({\mathbf{H}}_{E},{\mathcal{L}},f)-g)\|-\|\bm{\Phi}({\mathbf{H}}_{E},{\mathcal{L}},f)-g\|_{\mathcal{M}}\Big{|}\leq$
$\left(\frac{\log 1/\delta}{N}\right)^{\frac{1}{4}}\Bigg{)}\geq 1-\delta.$ (69)
Similar to the steps that we have in Appendix A-B, an expectation value can be derived based on the probability bound by decomposing the event space. We can have
$\mathbb{E}_{\mu^{N}}\left[\Big{|}\|{\mathbf{P}}_{N}(\bm{\Phi}({\mathbf{H}}_{E},{\mathcal{L}},f)-g)\|-\|\bm{\Phi}({\mathbf{H}}_{E},{\mathcal{L}},f)-g\|_{\mathcal{M}}\Big{|}\right]$
$\leq CN^{-1/4}+{\mathcal{O}}(e^{-N/C}\sqrt{N}).$ (70)
For the RHS in (65), we replace the loss function as $L_{2}$ loss and we have
$\Big{|}\ell(\bm{\Phi}({\mathbf{H}}_{E},{\mathbf{x}}_{N}),{\mathbf{y}}_{N})-\ell(\bm{\Phi}({\mathbf{H}}_{E},{\mathcal{L}},f),g)\Big{|}$
$\leq\Big{|}\|\bm{\Phi}({\mathbf{H}}_{E},{\mathbf{L}}_{N},{\mathbf{x}}_{N})-{\mathbf{P}}_{N}g\|-\|\bm{\Phi}({\mathbf{H}}_{E},{\mathcal{L}},f)-g\|_{\mathcal{M}}\Big{|}$ (71)
@@
By subtracting and adding an intermediate term ${\mathbf{P}}_{N}\bm{\Phi}({\mathbf{H}}_{E},{\mathcal{L}},f)$ as well as the triangle inequality we have the follows hold with probability at least $1-\delta$.
$\Big{|}\|\bm{\Phi}({\mathbf{H}}_{E},{\mathbf{L}}_{N},{\mathbf{x}}_{N})-{\mathbf{P}}_{N}g\|-\|\bm{\Phi}({\mathbf{H}}_{E},{\mathcal{L}},f)-g\|_{\mathcal{M}}\Big{|}$
$\leq\Big{|}\|\bm{\Phi}({\mathbf{H}}_{E},{\mathbf{L}}_{N},{\mathbf{x}}_{N})-{\mathbf{P}}_{N}\bm{\Phi}({\mathbf{H}}_{E},{\mathcal{L}},f)\|+$
$\quad\|{\mathbf{P}}_{N}\bm{\Phi}({\mathbf{H}}_{E},{\mathcal{L}},f)-{\mathbf{P}}_{N}g\|-\|\bm{\Phi}({\mathbf{H}}_{E},{\mathcal{L}},f)-g\|_{\mathcal{M}}\Big{|}$ (72)
$\leq\|\bm{\Phi}({\mathbf{H}}_{E},{\mathbf{L}}_{N},{\mathbf{x}}_{N})-{\mathbf{P}}_{N}\bm{\Phi}({\mathbf{H}}_{E},{\mathcal{L}},f)\|+$
$\Big{|}\|{\mathbf{P}}_{N}(\bm{\Phi}({\mathbf{H}}_{E},{\mathcal{L}},f)-g)\|-\|\bm{\Phi}({\mathbf{H}}_{E},{\mathcal{L}},f)-g\|_{\mathcal{M}}\Big{|}$ (73)
$\leq\|\bm{\Phi}({\mathbf{H}}_{E},{\mathbf{L}}_{N},{\mathbf{x}}_{N})-{\mathbf{P}}_{N}\bm{\Phi}({\mathbf{H}}_{E},{\mathcal{L}},f)\|+O\left(\left(\frac{\log(1/\delta)}{N}\right)^{\frac{1}{4}}\right).$ (74)
@@
Inequality (29) allows the derivation from (73) to (74). The first part in (74) is bounded by Proposition 1. Combining these two parts and taking the leading orders from (A-A), (70) and (60), we can achieve the conclusion that
$GA={\mathcal{O}}\left(\left(\frac{\log\frac{N}{\delta}}{N}\right)^{\frac{1}{d+4}}+\left(\frac{\log N}{N}\right)^{\frac{1}{d+4}}\right).$ (75)

### v2

@@ IV-B Cora Dataset
The Cora dataset is a smaller graph than the Arxiv dataset. The task is also node label prediction, and in this case, there are $7$ classes. In this dataset, the input feature dimension is $1,433$, with $2,708$ nodes and $5,429$ edges [46]. In this case the experiments were done for $\{16,32,64\}$ hidden units and $\{2,3\}$ layers.
@@
Visually, we verify that the generalization gap follows a linear decrease in log scale with respect to the number of nodes in the training set. This gap is driven by the increase in the testing accuracy/loss given that the training value remains constant. Formally, we can see in Table LABEL:table:pearson that the Pearson correlation coefficient is always above $0.96$ (except $2$ layers $64$ hidden units measured in the loss) which again shows a strong linear correlation between the number of nodes in the training set (in log scale) and the generalization gap.
@@
IV-C CiteSeer Dataset
The CiteSeer dataset has $3327$ nodes and $9228$ edges. There are $6$ classes, the feature dimension is $3703$, and we trained GNN with $\{16,32,64\}$ hidden units and $\{2,3\}$ layers.
In Figure 5 we can see the generalization gap results which show a linear correlation in log-scale. In Table LABEL:table:pearson we show the Pearson correlation coefficient for these experiments. For Accuracy, the results show a strong linear correlation as our theory predicts.
IV-D PubMed Dataset
The CiteSeer dataset has $19717$ nodes and $88651$ edges. There are $3$ classes and the feature dimension is $500$ dimensional, and we trained GNN with $\{16,32,64\}$ hidden units and $\{2,3\}$ layers.
In Figure 6 we can see the generalization gap results which show a linear correlation in log-scale. In Table LABEL:table:pearson we show the Pearson correlation coefficient for these experiments. Both for Accuracy and Loss, the results show a strong linear correlation as our theory predicts.
@@ V Conclusion
In this paper, we implemented a manifold convolutional neural network as a limit model for graph neural networks. We considered the case when graphs are constructed from samples from the manifold. We showed that when a graph is constructed with the i.i.d. randomly sampled points over the manifold, the GNN converges to the MNN with a non-asymptotic rate both in probability and in expectation. We utilized these convergence results as a stepping stone towards proving an upper bound on the generalization gap of GNNs. Formally, we showed that the difference between the statistical and empirical risks decreases with the number of sampled points from the manifold while it increases with the dimension of the manifold. Furthermore, our bound admits a linear interpretation in a logarithmic scale which can be validated empirically. Through experiments, we verified our convergence and generalization results numerically with two real-world datasets. We trained GNNs on a varying number of nodes and compared their generalization gaps. We empirically verified that the theoretically proposed rate matches the one seen in practice for a variety of layers and hidden dimensions. We also computed the Pearson Correlation coefficient to validate the linear relationships and for the most part, found values always above $0.9$, which indicates a strong linear relationship.
@@ A-A Proof of Proposition 1
[40, Theorem 2.4, Theorem 2.6] Let ${\mathcal{M}}\subset{\mathbb{R}}^{\mathsf{N}}$ of measure $\mu$ with probability density function $\rho$ be equipped with a weighted Laplace operator ${\mathcal{L}}_{\rho}$ as defined in (9), whose eigendecomposition is given by $\{\lambda_{i},\bm{\phi}_{i}\}_{i=1}^{\infty}$. Let ${\mathbf{L}}_{N}$ be the discrete graph Laplacian of graph weights defined as (15), with spectrum $\{\lambda_{i,N},\bm{\phi}_{i,N}\}_{i=1}^{N}$. Fix $K\in\mathbb{N}^{+}$ and assume that $\epsilon=\epsilon(N)\geq\left({\log(CN/\delta)}/{N}\right)^{1/(d+4)}$ Then, with probability at least $1-\delta$, we have
@@
$\leq\sum_{i=M+1}^{N}\left|\hat{h}(\lambda_{i,N})\right|\|{\mathbf{P}}_{N}f\|\|\bm{\phi}_{i,N}\|$ (46)
@@ A-C Proof of Theorem 1
$\min_{{\mathbf{H}}\in{\mathcal{H}}}R_{S}({\mathbf{H}})\leq R_{S}({\mathbf{H}}_{E}),\quad\min_{{\mathbf{H}}\in{\mathcal{H}}}R_{E}({\mathbf{H}})=R_{E}({\mathbf{H}}_{E}),$ (62)
which leads to $GA=\min_{{\mathbf{H}}\in{\mathcal{H}}}R_{S}({\mathbf{H}})-R_{E}({\mathbf{H}}_{E})$ bounded as
$GA\leq R_{S}({\mathbf{H}}_{E})-R_{E}({\mathbf{H}}_{E})$ (63)
$=\mathbb{E}_{X_{N}}\left[\ell\left(\bm{\Phi}({\mathbf{H}}_{E},{\mathbf{L}}_{N},{\mathbf{x}}_{N}),{\mathbf{y}}_{N}\right)\right]-\ell(\bm{\Phi}({\mathbf{H}}_{E},{\mathbf{L}}_{N},{\mathbf{x}}_{N}),{\mathbf{y}}_{N}).$ (64)
We can now subtract and add an intermediate-term $\mathbb{E}[\ell({\mathbf{P}}_{N}\bm{\Phi}({\mathbf{H}}_{E},{\mathcal{L}},f),{\mathbf{P}}_{N}g)]$, where the expectation is taken over all the random sampling operator ${\mathbf{P}}_{N}$ that results in the random sampling point set $X_{N}$. The $GA$ term therefore can be bounded by
$GA\leq\Bigg{(}\mathbb{E}_{X_{N}}\left[\ell\left(\bm{\Phi}({\mathbf{H}}_{E},{\mathbf{L}}_{N},{\mathbf{x}}_{N}),{\mathbf{y}}_{N}\right)\right]$
$\qquad\qquad\qquad\qquad\qquad-\mathbb{E}[\ell({\mathbf{P}}_{N}\bm{\Phi}({\mathbf{H}}_{E},{\mathcal{L}},f),{\mathbf{P}}_{N}g)]\Bigg{)}+$
$\Bigg{(}\mathbb{E}[\ell({\mathbf{P}}_{N}\bm{\Phi}({\mathbf{H}}_{E},{\mathcal{L}},f),{\mathbf{P}}_{N}g)]-\ell(\bm{\Phi}({\mathbf{H}}_{E},{\mathbf{L}}_{N},{\mathbf{x}}_{N}),{\mathbf{y}}_{N})\Bigg{)}.$ (65)
@@
$GA\leq\Bigg{|}\mathbb{E}_{X_{N}}\left[\ell\left(\bm{\Phi}({\mathbf{H}}_{E},{\mathbf{L}}_{N},{\mathbf{x}}_{N}),{\mathbf{y}}_{N}\right)\right]$
$\qquad\qquad\qquad\qquad\qquad-\mathbb{E}[\ell({\mathbf{P}}_{N}\bm{\Phi}({\mathbf{H}}_{E},{\mathcal{L}},f),{\mathbf{P}}_{N}g)]\Bigg{|}+$
$\left|\mathbb{E}[\ell({\mathbf{P}}_{N}\bm{\Phi}({\mathbf{H}}_{E},{\mathcal{L}},f),{\mathbf{P}}_{N}g)]-\ell(\bm{\Phi}({\mathbf{H}}_{E},{\mathbf{L}}_{N},{\mathbf{x}}_{N}),{\mathbf{y}}_{N})\right|.$ (66)
The first term in (66) can be analyzed by combining the expectation as the random sampling operator leads to the random sampling set.
$\left|\mathbb{E}_{X_{N}}\left[\ell\left(\bm{\Phi}({\mathbf{H}}_{E},{\mathbf{L}}_{N},{\mathbf{x}}_{N}),{\mathbf{y}}_{N}\right)\right]-\mathbb{E}[\ell({\mathbf{P}}_{N}\bm{\Phi}({\mathbf{H}}_{E},{\mathcal{L}},f),{\mathbf{y}}_{N})]\right|$
$=\left|\mathbb{E}[\ell\left(\bm{\Phi}({\mathbf{H}}_{E},{\mathbf{L}}_{N},{\mathbf{x}}_{N}),{\mathbf{y}}_{N}\right)-\ell({\mathbf{P}}_{N}\bm{\Phi}({\mathbf{H}}_{E},{\mathcal{L}},f),{\mathbf{y}}_{N})]\right|$ (67)
$\leq\mathbb{E}\left[\left|\ell\left(\bm{\Phi}({\mathbf{H}}_{E},{\mathbf{L}}_{N},{\mathbf{x}}_{N}),{\mathbf{y}}_{N}\right)-\ell({\mathbf{P}}_{N}\bm{\Phi}({\mathbf{H}}_{E},{\mathcal{L}},f),{\mathbf{y}}_{N})\right|\right]$ (68)
Considering the Lipschitz continuity assumption of the loss function in Assumption 2, we have
$\mathbb{E}\left[\left|\ell\left(\bm{\Phi}({\mathbf{H}}_{E},{\mathbf{L}}_{N},{\mathbf{x}}_{N}),{\mathbf{y}}_{N}\right)-\ell({\mathbf{P}}_{N}\bm{\Phi}({\mathbf{H}}_{E},{\mathcal{L}},f),{\mathbf{y}}_{N})\right|\right]$ (69)
$\leq\mathbb{E}\left[\|\bm{\Phi}({\mathbf{H}}_{E},{\mathbf{L}}_{N},{\mathbf{x}}_{N})-{\mathbf{P}}_{N}\bm{\Phi}({\mathbf{H}}_{E},{\mathcal{L}},f),{\mathbf{y}}_{N})\|_{2}\right],$ (70)
which can be bounded by the conclusion in Corollary 2.
@@
For the second term in (66), we subtract and add an intermediate-term $\ell({\mathbf{P}}_{N}\bm{\Phi}({\mathbf{H}}_{E},{\mathcal{L}},f),{\mathbf{P}}_{N}g)$, with the triangle inequality, we have
$\left|\mathbb{E}[\ell({\mathbf{P}}_{N}\bm{\Phi}({\mathbf{H}}_{E},{\mathcal{L}},f),{\mathbf{P}}_{N}g)]-\ell(\bm{\Phi}({\mathbf{H}}_{E},{\mathbf{L}}_{N},{\mathbf{x}}_{N}),{\mathbf{y}}_{N})\right|$
$\leq\left|\mathbb{E}[\ell({\mathbf{P}}_{N}\bm{\Phi}({\mathbf{H}}_{E},{\mathcal{L}},f),{\mathbf{P}}_{N}g)]-\ell({\mathbf{P}}_{N}\bm{\Phi}({\mathbf{H}}_{E},{\mathcal{L}},f),{\mathbf{P}}_{N}g)\right|$
$+\left|\ell({\mathbf{P}}_{N}\bm{\Phi}({\mathbf{H}}_{E},{\mathcal{L}},f),{\mathbf{P}}_{N}g)-\ell(\bm{\Phi}({\mathbf{H}}_{E},{\mathbf{L}}_{N},{\mathbf{x}}_{N}),{\mathbf{y}}_{N})\right|.$ (71)
The first term in (71) is the deviation of the random term $\ell({\mathbf{P}}_{N}\bm{\Phi}({\mathbf{H}}_{E},{\mathcal{L}},f),{\mathbf{P}}_{N}g)$, which can be bounded by Chebyshev’s inequality as both the output value and the target functions are bounded. The second term in (71) can be bounded based on Assumption 2 and Proposition 1 as
$\left|\ell({\mathbf{P}}_{N}\bm{\Phi}({\mathbf{H}}_{E},{\mathcal{L}},f),{\mathbf{P}}_{N}g)-\ell(\bm{\Phi}({\mathbf{H}}_{E},{\mathbf{L}}_{N},{\mathbf{x}}_{N}),{\mathbf{y}}_{N})\right|$
$\leq\|{\mathbf{P}}_{N}\bm{\Phi}({\mathbf{H}}_{E},{\mathcal{L}},f)-\bm{\Phi}({\mathbf{H}}_{E},{\mathbf{L}}_{N},{\mathbf{x}}_{N})\|_{2}$ (72)
@@
Combining the upper bounds of these two terms and taking the leading orders from (A-A) and (60), we can achieve the conclusion that
$GA={\mathcal{O}}\left(\left(\frac{\log\frac{N}{\delta}}{N}\right)^{\frac{1}{d+4}}\right).$ (73)